Can you recognise those different parts of trigonometry, those basic ideas finding the different sides and labeling them. Can we recognise it in daily life? Find the right side of the triangle.labels the sides or specific angle that you find.
1
Expert's answer
2020-04-08T13:29:06-0400



From the above figure we can see that there is a right angled triangle then the side opposite to the 90o angle is hypotenuse , and side adjacent to the angle "\theta" is called base and the side opposite to the angle "\theta" is called height.

From the above explanation we can see that ABC is a right angled triangle, right angle at B so side opposite is called hypotenuse which is AC , we can see that "\angle BAC=\theta" so the side opposite to the "\theta" is called height which is BC here and the side adjacent to the "\theta" is called base which is AB here.


"\sin\theta=\frac{height}{hypotenuse}=\frac{BC}{AC}"



"\cos\theta=\frac{base}{hypotenuse}=\frac{AB}{AC}"


"\tan\theta=\frac{height}{base }=\frac{BC}{AB}"